# Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### 17 Jan 2024 v3

@@ 4.2 TTM’s Zero/Few-shot Performance
Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings across multiple $fl$s. Baseline results are reported from  Zhou et al. (2023) as we use the same few-shot data filtering strategy as followed in that paper. Note that for zero-shot performance, the pre-trained TTM model is directly evaluated on the test set. The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%). An even more impressive observation is that the TTM in zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data. This observation establishes the generalization ability of the pre-trained TTM model on the target datasets. Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains. In addition, TTM zero-shot also beats many SOTAs (not all) with 10% training highlighting the effectiveness of our approach.
Additionally, we conduct a comparison between TTM and the LLMTime model which is explicitly designed for zero-shot setting. Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set . Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  Gruver et al. (2023). In this evaluation, we outperform LLMTime by a substantial margin of 29%. In addition, there are alternative pre-training approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected. Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al. (2023)). Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%. Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs. Notably, we achieve these accuracy gains with significantly reduced compute requirements, as elaborated next.
@@ 4.4 TTM’s effectiveness in cross-channel and exogenous modeling
Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits. Since these datasets are already very small, we used their full data for fine-tuning. Table 5 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare the performance of TTM-CM with the pre-trained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Other SOTAs are not reported here, considering their inferior performance and space constraints. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al. (2023) as they are the latest SOTAs in channel-correlation modeling. We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
@@ 4.5 Ablation Studies
Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone. Table 7 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module. We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot. Performance improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.

### 09 Apr 2024 v5

@@ 4.2 TTM’s Zero/Few-shot Performance
Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings across multiple $fl$s. Baseline results are reported from  [?] as we use the same few-shot data filtering strategy as followed in that paper. Note that for zero-shot performance, the pre-trained TTM model is directly evaluated on the test set. The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%). An even more impressive observation is that the TTM in zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data. This observation establishes the generalization ability of the pre-trained TTM model on the target datasets. Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains. In addition, TTM zero-shot also beats many SOTAs (not all) with 10% training highlighting the effectiveness of our approach.
Additionally, we conduct a comparison between TTM and the LLMTime model which is explicitly designed for zero-shot setting. Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set . Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  [?]. In this evaluation, we outperform LLMTime by a substantial margin of 29%. In addition, there are alternative pre-training approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected. Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  [?]). Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%. Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs. Notably, we achieve these accuracy gains with significantly reduced compute requirements, as elaborated next.
@@ 4.4 TTM’s effectiveness in cross-channel and exogenous modeling
Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits. Since these datasets are already very small, we used their full data for fine-tuning. Table 5 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare the performance of TTM-CM with the pre-trained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Other SOTAs are not reported here, considering their inferior performance and space constraints. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) [?] as they are the latest SOTAs in channel-correlation modeling. We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
@@ 4.5 Ablation Studies
Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1 is a key component of the TTM backbone. Table 4.2 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module. We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot. Performance improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
@@
4.6 Comparison with recent concurrent works
In this section, we compare the performance of TTM with the very recent concurrent works published in 2024. Time-LLM [?] is a recent SOTA that enables successful reprogramming of LLMs for time-series tasks. However, as depicted in Table 10, TTM outperforms Time-LLM by 8% in Fewshot 5% and 2% in Fewshot 10% setting. Likewise, UniTime [?] proposes novel techniques towards language-empowered cross-domain time-series forecasting. As indicated in Table 11, TTM’s zero-shot results surpass UniTime’s zero-shot results by 27%. Further, TTM does not need any cross-domain transfer learning and pair-wise mapping as our pretraining is enabled on a diverse set of datasets.
Data Few-shot (5%)) Few-shot (10%))
TTM Time-LLM TTM Time-LLM
ETTH1 0.393 0.627 0.422 0.556
ETTH2 0.334 0.382 0.36 0.37
ETTM1 0.44 0.425 0.458 0.404
ETTM2 0.28 0.274 0.278 0.277
Weather 0.231 0.26 0.227 0.234
Electricity 0.179 0.179 0.18 0.175
Traffic 0.421 0.423 0.428 0.429
TTM IMP(%) 8% 2%
Table 10: TTM Vs Time-LLM on D1 Datasets averaged across considered 4 forecast lengths. MSE reported.
Data $fl$ Zero-shot
UniTime TTM
ETTH2 96 0.306 0.285
192 0.389 0.341
336 0.424 0.383
720 0.434 0.441
Electricity 96 0.409 0.169
192 0.41 0.196
336 0.439 0.209
720 0.487 0.264
Weather 96 0.21 0.154
192 0.264 0.203
336 0.326 0.256
720 0.402 0.329
TTM IMP (%) 27%
Table 11: TTM vs UniTime
MSE Improvement (IMP)
in zero-shot setting.